OpenStudy (anonymous):

Solve the equation x-2/x+4 + x+4/x-2=17/4 . . . . ANSWER = {4,-6}

OpenStudy (anonymous):

take LCM

OpenStudy (anonymous):

\[\frac{ x-2 }{ x+4 }+\frac{ x+4 }{ x-2 }=\frac{ 17 }{ 4}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what you got on taking LCM , and then simplifying.

OpenStudy (anonymous):

\[\frac{ 2x^2+4x+20 }{ x^2+2x-8 }=\frac{ 17 }{ 4 }\]

OpenStudy (anonymous):

now cross multiply

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what eq. you got?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

\[9x^2+18x-216=0\]

OpenStudy (anonymous):

divide by 9

OpenStudy (anonymous):

then factorise

OpenStudy (anonymous):

how will it factorize

OpenStudy (anonymous):

\[x^{2}+2x-24=0\] \[x^{2}+6x-2x-24=0\]

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

second equation not equal first equation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

something went wrong

OpenStudy (anonymous):

soryy

OpenStudy (anonymous):

it is 6x-4x

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

x=-6 and x=4